Find integer such that x + y = 5, x - y = 3
make sure that you prove that you have found all the solutions
thank you

Answer by CubeyThePenguin(3113)

x + y = 5
x - y = 3

Add the equations: 2x = 8 ---> x = 4

Substitute the result into any of the equations.

x + y = 5
4 + y = 5
y = 1

The solution is (x, y) = (4, 1).

To check:

4 + 1 = 5
4 - 1 = 3